OpenStudy (anonymous):

Please help figure out the function rule. Medal Awarded: x,y = 1,7; 2,29; 3,133; 4,641; 5,3157. Thanks.

OpenStudy (joannablackwelder):

I haven't finished it yet, but it looks like it will probably be some kind of power.

OpenStudy (anonymous):

They said it was expontential

OpenStudy (joannablackwelder):

3*1^3 +5 = 7 3*2^3 + 5 = 29

OpenStudy (joannablackwelder):

But that doesn't work for x=3

OpenStudy (joannablackwelder):

But I think it is something similar

OpenStudy (anonymous):

Here are the answers they said it was: a^x + b^x = y

OpenStudy (joannablackwelder):

Oh, ok.

OpenStudy (anonymous):

I have no idea..... : 0+7, 1+6, 2+5, 3+4,

OpenStudy (joannablackwelder):

Using the form you gave, we need to pick numbers for a and b, that when we plug in 1 for x, we get 7 for y.

OpenStudy (joannablackwelder):

If we make a=3 and b=4, the formula works when x=1\[3^{1}+4^{1}=7\]

OpenStudy (joannablackwelder):

But that form doesn't work for x=2\[3^{2}+4^{2}=25\]

OpenStudy (joannablackwelder):

We needed it to be 29, so we need to pick difference numbers for a and b.

OpenStudy (joannablackwelder):

Does that make sense so far?